OpenStudy (anonymous):

David weights 5kg less than John, who in turn weighs 8kg less than Paul, their total weight is 197kg, find the weight of each person. I wrote paul's weight =x so David = x-5 and John = x-8 = x-5+x-8=197???

OpenStudy (anonymous):

let paul=x john=x-8 david=x-8-5=x-13 so, 197=x+(x-8)+(x-13)

OpenStudy (anonymous):

because john less than david

OpenStudy (anonymous):

so need to minus 8 more kg to david

OpenStudy (anonymous):

and "total" mean add up THREE of them not TWO fo them

OpenStudy (anonymous):

do you remember the other word problem you solved for me?

OpenStudy (anonymous):

ya.. any question??

OpenStudy (anonymous):

they seemed the same to me, thats why I tried to solve it exactly the same way as you showed me in the the first question I posted.